OpenStudy (anonymous):

x+y=5 and xy=15, find x^2+y^2

OpenStudy (anonymous):

Use: \[\large x^2 + y^2 = (x+y)^2 - 2xy\]

OpenStudy (anonymous):

Solve this: \[\large x^2 + y^2 = (5)^2 - 2(15)\] Can you do it now??

OpenStudy (anonymous):

yes.... thank you

OpenStudy (anonymous):

Wait how the result can be negative?? I am in doubt now.. Wait..

OpenStudy (anonymous):

k

OpenStudy (klimenkov):

\(x_1=\frac12(5-i\sqrt{35})\) \(x_2=\frac12(5+i\sqrt{35})\) \(x_1,x_2\) are complex numbers, so it is okay.